# SemCoT: Accelerating Chain-of-Thought Reasoning through Semantically-Aligned Implicit Tokens

### v1

@@ 1 Introduction
Chain-of-Thought (CoT) wei2022chain is a technique where Large Language Models (LLMs) demonstrate step-by-step reasoning by breaking down complex questions into sequential steps. This approach can be implemented through finetuning with specialized datasets liuchain ; yu2023alert or providing crafted prompt instructions wei2022chain . CoT has achieved remarkable performance in various reasoning-intensive NLP tasks such as mathematical problem solving jie2023design ; miner2024scheherazade ; lyu2023faithful and symbolic reasoning xu2024faithful ; lyu2023faithful ; sprague2024cot . In light of the impressive performance gains achieved through Chain-of-Thought (CoT) prompting, several advanced reasoning models have recently been developed, including OpenAI’s o1/o3 models jaech2024openai ; OpenAI2025 and DeepSeek R1 guo2025deepseek . These models have been specifically trained to tackle problems that require complex, long-chain reasoning. However, as CoT reasoning becomes particularly verbose for LLMs with extensive parameters, it also significantly extends LLMs’ reasoning time. For example, ChatGPT-4o hurst2024gpt may generate as many as around five hundred tokens during CoT reasoning (taking up to 21.37 seconds cheng2024compressed ) to complete a grade school-level math problem, whose answer is simply a numerical number with two to four tokens. Several works aim to improve CoT efficiency, with one primary strategy being implicit CoT deng2024explicit ; hao2024training ; shen2025efficient ; cheng2024compressed ; su2025token ; xu2025softcot ; shen2025codi ; deng2023implicit , as shown in Fig. 1  [^1] The figure shows how most (not all) implicit CoT methods perform inference with a query.. Implicit CoT substitutes the explicit reasoning tokens with a small number of LLM’s first-layer hidden embeddings (each embedding is called an implicit CoT token). Implicit CoT can improve efficiency because it not only pursues more concise reasoning (box in the middle of Fig. 1 (b)) but also avoids going through unembedding layers (box at the upper side in Fig. 1 (a)) to decode from last-layer hidden embeddings to tokens, and the tokenization (box at the lower side in Fig. 1 (a)) of reasoning tokens.
@@
However, current implicit CoT methods encounter two significant challenges: (1) Gap towards Maintaining Effectiveness: Current implicit CoT methods struggle to establish semantic alignment between implicit embeddings (when transformed to natural language) and ground-truth reasoning. This “semantic alignment difficulty” stems from a fundamental format mismatch: implicit CoT exists as hidden embeddings while ground-truth reasoning consists of natural language. Existing approaches inadequately address this by either: (i) completely discarding ground-truth reasoning goyalthink ; xu2025softcot , (ii) matching only keywords from ground-truth reasoning cheng2024compressed , or (iii) first finetune LLMs with ground-truth reasoning, then optimizing implicit reasoning only for generating correct answers hao2024training ; deng2024explicit . (2) Gap towards Enhancing Efficiency: Although some implicit CoT methods cheng2024compressed ; shen2025codi ; hao2024training ; deng2023implicit reduce reasoning length, they neglect the high computational cost for generating each reasoning token with the LLM. This challenge becomes particularly important when LLMs scale to substantial sizes (hundreds of billions of parameters) nowadays. For example, generating one token can cost as high as approximately 0.1s for DeepSeek-R1 deepseekai2025deepseekr1 ; parisi2025deepseekr1 . The computational overhead accumulates significantly during reasoning, especially for complex problems requiring extensive step-by-step thinking.
@@
To tackle the above challenges, in this paper, we propose a novel framework named SemCoT (Semantically-aligned Implicit CoT) to accelerate implicit CoT while effectively preserving the performance benefits of traditional CoT methods. Specifically, to tackle the first challenge, we propose a customized sentence transformer to measure the semantic alignment between implicit reasoning and ground-truth reasoning. The customized sentence transformer converts each of the implicit reasoning and ground-truth reasoning into an embedding vector that characterizes its semantics. Thus, we can compare their semantic alignment by cosine similarity chandrasekaran2021evolution . The well-trained, customized sentence transformer is utilized to optimize implicit CoT generation to be semantically aligned with ground-truth reasoning. To tackle the second challenge, we adopt a lightweight language model (LM)  [^2] As the lightweight model may not be a large language model, we only call it a language model., which is an off-the-shelf LM that is distilled or sheared (i.e., pruned) from the original LLM, to generate implicit reasoning. This approach dramatically reduces the time cost of generating each single CoT token. The lightweight LM is guided by our sentence transformer to distill ground-truth reasoning into semantically aligned implicit reasoning, while also optimized for answer accuracy gou2021knowledge .
@@ 3.1 Overview
Motivation. As introduced in Section 1, existing methods struggle to fully capture ground-truth reasoning information in implicit reasoning. We address this challenge with a natural assumption: optimal implicit CoT performance 333“Implicit CoT performance” refers to the LLM answer accuracy with the implicit CoT reasoning. is achieved when the implicit CoT semantically aligns with the ground-truth CoT. This alignment occurs when the implicit CoT, if translated from the LLM’s embedding space to human language, would be semantically equivalent to the ground-truth reasoning. However, this “translation” presents a fundamental challenge: we cannot formulate it in explicit formulations due to its complexity, nor can we model it with neural networks because human language space is discrete and therefore not differentiable. As a practical alternative, we measure this alignment within the LLM’s embedding space by comparing the input embedding of ground-truth reasoning (what we call “embedded ground-truth reasoning”) with the implicit reasoning. Although one might consider directly using vector distance metrics to measure semantic alignment between implicit reasoning and embedded ground-truth reasoning, such approaches have significant limitations: (1) LLM embeddings are optimized for next-token prediction rather than sentence semantics, vector distance may not faithfully reveal semantic similarity between sentencesreimers2019sentence . (2) LLM embeddings are high dimensional and suffer from the curse of dimensionality verleysen2005curse , where distance metrics become less meaningful as dimensionality increases huang2023advancing .
@@
Customized Sentence Transformer. Recent advances in sentence transformers reimers2019sentence ; cer2018universal ; gao2021simcse offer state-of-the-art methods to measure semantic relationships between sentences. They utilize a transformer-based architecture combined with pooling and linear layers to produce semantic vector representations of input sentences, enabling the measurement of sentence similarity via cosine similarity. Although traditional sentence transformers cannot be directly applied in our context—since implicit reasoning steps are not explicitly formulated as sentences—they inspire our approach. A naive solution would involve bypassing the tokenization step of sentence transformers and directly feeding implicit or embedded ground-truth reasoning into these models. However, this approach is inadequate, as standard sentence transformers and the LLM we examine typically employ different mappings from tokens to input embeddings, resulting in embedding spaces that differ semantically. To address this issue, we design a customized sentence transformer specifically tailored to handle embeddings from LLMs, thereby accurately measuring the semantic alignment between reasoning.
@@
We design a customized sentence transformer, shown in Fig. 2 (a), denoted as $\mathcal{C}_{\phi}$ (with parameters $\phi$), tailored to comparing semantics between implicit reasoning and ground-truth reasoning. First, we extract the middle five layers of the LLM that we perform CoT on to serve as the backbone of our customized sentence transformer, as these layers have been shown in prior work liu2019linguistic to possess optimal language modeling ability and transferability across tasks. Next, we add a pooling layer on top of the transformer layers to aggregate token embeddings from the entire reasoning sequence into a unified vector representation. Finally, we employ a linear layer to project this unified vector into a lower-dimensional semantic embedding space, facilitating more efficient similarity comparisons.
@@
The customized sentence transformer is trained with a specially crafted reasoning pair dataset $\mathcal{G}$, where each data point $(R,S)\in\mathcal{G}$ is a reasoning pair with $R$ being the ground-truth reasoning and $S$ a GPT-4o-mini OpenAI2024GPT4omini generated condensed semantically aligned reasoning of $R$. The prompt to generate $S$ with GPT-4o-mini OpenAI2024GPT4omini is “Please generate the most condensed reasoning text which is semantically aligned with the following reasoning text: RR.” We train the sentence transformer with a contrastive learning le2020contrastive strategy. Specifically, we compute the semantic embeddings of the ground-truth reasoning $\mathcal{C}_{\phi}(\mathcal{T}_{\mathcal{F}}(R_{i}))$ and condensed reasoning $\mathcal{C}_{\phi}(\mathcal{T}_{\mathcal{F}}(S_{i}))$, where $\mathcal{T}_{\mathcal{F}}$ is the mapping from tokens to input embeddings of the LLM. The sentence transformer is trained using contrastive learning to maximize the similarity between embeddings of positive pairs $(R_{i},S_{i})$, while minimizing similarities between all negative pairs $(R_{i},S_{j})$ for $j\neq i$ within the same minibatch. Formally, we learn the sentence transformer with
@@ 3.3 Token-level Efficient Implicit Reasoning Generation
Lightweight Implicit Reasoning Generator. We build the lightweight implicit reasoning generator, shown on the right in Fig. 2 (a), based on an off-the-shelf pruned or distilled LM (e.g., Sheared-LLaMA-1.3B xiasheared ) from the LLM (e.g., Llama-2-7b-chat-hf llama2 ). We choose to use a pruned or distilled LMs of the original LLMs because existing research tao2024llms ; wang2025multi suggests that those models preserve crucial semantic properties of their original LLMs in their embedding spaces. Then, we incorporate a linear projection layer to map the last-layer hidden embedding generated by the lightweight LM into the embedding space of the LLM to serve as implicit reasoning for the LLM.
@@ 4.1 Experiment Settings
Datasets. We adopt five representative datasets from three different semantic domains used for benchmarking CoT performance. Specifically, we apply mathematical reasoning datasets GSM8K cobbe2021training , SVAMP patel-etal-2021-nlp , MultiArith ChilleD2023MultiArith , commonsense reasoning dataset CommonsenseQA talmor-etal-2019-commonsenseqa , and symbolic reasoning dataset CoinFlip krishna2023coinflip . Please see Appendix C for the metadata of the five used datasets.
Examined LLM & Implicit Reasoning Generator. We examine SemCoT on two representative open source LLMs, Llama-2-7b-chat-hf llama2 and Mistral-7B-Instruct-v0.2 mistral . For the lightweight LM within the implicit reasoning generator, we utilize the distilled/sheared LMs from their corresponding examined LLMs. Specifically, in the cases that the LLM is Llama-2-7b-chat-hf llama2 , we employ Sheared-LLaMA-1.3B xiasheared as the lightweight LM within the implicit reasoning generator. When the backbone LLM is Mistral-7B-Instruct-v0.2 mistral , we apply mistral-1.1b-testing optimum2024mistral accordingly.
Baselines. We adopt state-of-the-art baselines improving CoT efficiency on LLMs. Specifically, (1) Pause goyalthink uses identical implicit reasoning tokens to substitute ground-truth reasoning. Their training strategy includes pretraining and finetuning the LLM for optimal answer accuracy with implicit reasoning. We only implement the finetuning strategy due to the significant cost of LLM pretraining. (2) Progressive encoding methods, such as ICoT-SI deng2024explicit and COCONUT xu2025softcot , use implicit tokens to substitute explicit tokens gradually during LLM finetuning. (3) CODI shen2025codi adopts a self-distillation strategy, where the teacher and student are both the LLM. The teacher learns to perform explicit reasoning, while the student learns to generate implicit tokens that lead to the correct answer. Simultaneously, both models are trained to align their token embeddings at the final token position of the query. (4) SoftCoT xu2025softcot utilizes a small LM to generate LLM’s hidden reasoning tokens without compressing the number of hidden tokens compared with the original explicit reasoning chain.
@@
Implementation Details. We set the output embedding dimension of the customized sentence transformer to 768. The number of implicit tokens during training is five, and, during evaluation, it is set to one. We optimize both the customized sentence transformer and the implicit reasoning generator with AdamW loshchilovdecoupled using the best hyperparameters found. For inference, we allow up to thirty answer tokens to be generated to enforce the LLM to generate the answer instead of the CoT.
@@ 4.2 Effectiveness & Efficiency of SemCoT
In this subsection, we aim to answer RQ1. Specifically, we evaluate our proposed framework SemCoT on two LLMs, Llama-2-7b-chat-hf llama2 and Mistral-7B-Instruct-v0.2 mistral . We compare the two LLMs’ answer accuracy and inference time cost when performing CoT with our SemCoT and the state-of-the-art baselines using one implicit reasoning token. We show the results in Table 1. From Table 1, we can make the following observations: (1) from the perspective of effectiveness, our SemCoT achieves the highest answer accuracy compared with the baseline methods in most datasets, showing that our method effectively retains LLM’s reasoning ability in a wide spectrum of NLP tasks and LLM settings. (2) From the perspective of efficiency, SemCoT achieves nearly the fastest implicit reasoning processing time across all datasets and LLMs. In some datasets, SemCoT is slightly slower than SoftCoT, but SemCoT achieves much higher answer accuracy in those cases. In conclusion, SemCoT achieves the optimal performance w.r.t. the trade-off between efficiency and effectiveness.

### v2

@@ 1 Introduction
Chain-of-Thought (CoT) Wei et al. (2022) is a technique where Large Language Models (LLMs) demonstrate step-by-step reasoning by breaking down complex questions into sequential steps. This approach can be implemented through finetuning with specialized datasets Liu et al. (2024); Yu et al. (2023) or providing crafted prompt instructions Wei et al. (2022). CoT has achieved remarkable performance in various reasoning-intensive NLP tasks such as mathematical problem solving Jie et al. (2023); Miner et al. (2024); Lyu et al. (2023) and symbolic reasoning Xu et al. (2024); Lyu et al. (2023); Sprague et al. (2025). In light of the impressive performance gains achieved through Chain-of-Thought (CoT) prompting, several advanced reasoning models have recently been developed, including OpenAI’s o1/o3 models Jaech et al. (2024); Zhang et al. and DeepSeek R1 Guo et al. (2025). These models have been specifically trained to tackle problems that require complex, long-chain reasoning. However, as CoT reasoning becomes particularly verbose for LLMs with extensive parameters, it also significantly extends LLMs’ reasoning time. For example, ChatGPT-4o Hurst et al. (2024) may generate as many as around five hundred tokens during CoT reasoning (taking up to 21.37 seconds Cheng and Van Durme (2024)) to complete a grade school-level math problem, whose answer is simply a numerical number with two to four tokens. Several works aim to improve CoT efficiency, with one primary strategy being implicit CoT Deng et al. (2024); Hao et al. (2024); Shen et al. (2025a); Cheng and Van Durme (2024); Su et al. (2025); Xu et al. (2025); Shen et al. (2025b); Deng et al. (2023), as shown in Fig. 1  [^1] The figure shows how most (not all) implicit CoT methods perform inference with a query.. Implicit CoT substitutes the explicit reasoning tokens with a small number of LLM’s first-layer hidden embeddings (each embedding is called an implicit CoT token). Implicit CoT can improve efficiency because it not only pursues more concise reasoning (box in the middle of Fig. 1 (b)) but also avoids going through unembedding layers (box at the upper side in Fig. 1 (a)) to decode from last-layer hidden embeddings to tokens, and the tokenization (box at the lower side in Fig. 1 (a)) of reasoning tokens.
@@
However, current implicit CoT methods encounter two significant challenges: (1) Gap towards Maintaining Effectiveness: Current implicit CoT methods struggle to establish semantic alignment between implicit embeddings (when transformed to natural language) and ground-truth reasoning. This “semantic alignment difficulty” stems from a fundamental format mismatch: implicit CoT exists as hidden embeddings while ground-truth reasoning consists of natural language. Existing approaches inadequately address this by either: (i) completely discarding ground-truth reasoning Goyal et al. (2024); Xu et al. (2025), (ii) matching only keywords from ground-truth reasoning Cheng and Van Durme (2024), or (iii) first finetune LLMs with ground-truth reasoning, then optimizing implicit reasoning only for generating correct answers Hao et al. (2024); Deng et al. (2024). (2) Gap towards Enhancing Efficiency: Although some implicit CoT methods Cheng and Van Durme (2024); Shen et al. (2025b); Hao et al. (2024); Deng et al. (2023) reduce reasoning length, they neglect the high computational cost for generating each reasoning token with the LLM. This challenge becomes particularly important when LLMs scale to substantial sizes (hundreds of billions of parameters) nowadays. For example, generating one token can cost as high as approximately 0.1s for DeepSeek-R1 Guo et al. (2025); Parisi (2025). The computational overhead accumulates significantly during reasoning, especially for complex problems requiring extensive step-by-step thinking.
@@
To tackle the above challenges, in this paper, we propose a novel framework named SemCoT (Semantically-aligned Implicit CoT) to accelerate implicit CoT while effectively preserving the performance benefits of traditional CoT methods. Specifically, to tackle the first challenge, we propose a customized sentence transformer to measure the semantic alignment between implicit reasoning and ground-truth reasoning. The customized sentence transformer converts each of the implicit reasoning and ground-truth reasoning into an embedding vector that characterizes its semantics. Thus, we can compare their semantic alignment by cosine similarity Chandrasekaran and Mago (2021). The well-trained, customized sentence transformer is utilized to optimize implicit CoT generation to be semantically aligned with ground-truth reasoning. To tackle the second challenge, we adopt a lightweight language model (LM)  [^2] As the lightweight model may not be a large language model, we only call it a language model., which is an off-the-shelf LM that is distilled or sheared (i.e., pruned) from the original LLM, to generate implicit reasoning. This approach dramatically reduces the time cost of generating each single CoT token. The lightweight LM is guided by our sentence transformer to distill ground-truth reasoning into semantically aligned implicit reasoning, while also optimized for answer accuracy Gou et al. (2021).
@@ 3.1 Overview
Motivation. As introduced in Section 1, existing methods struggle to fully capture ground-truth reasoning information in implicit reasoning. We address this challenge with a natural assumption: optimal implicit CoT performance 333“Implicit CoT performance” refers to the LLM answer accuracy with the implicit CoT reasoning. is achieved when the implicit CoT semantically aligns with the ground-truth CoT. This alignment occurs when the implicit CoT, if translated from the LLM’s embedding space to human language, would be semantically equivalent to the ground-truth reasoning. However, this “translation” presents a fundamental challenge: we cannot formulate it in explicit formulations due to its complexity, nor can we model it with neural networks because human language space is discrete and therefore not differentiable. As a practical alternative, we measure this alignment within the LLM’s embedding space by comparing the input embedding of ground-truth reasoning (what we call “embedded ground-truth reasoning”) with the implicit reasoning. Although one might consider directly using vector distance metrics to measure semantic alignment between implicit reasoning and embedded ground-truth reasoning, such approaches have significant limitations: (1) LLM embeddings are optimized for next-token prediction rather than sentence semantics, vector distance may not faithfully reveal semantic similarity between sentencesReimers and Gurevych (2019). (2) LLM embeddings are high dimensional and suffer from the curse of dimensionality Verleysen and François (2005), where distance metrics become less meaningful as dimensionality increases Huang et al. (2023).
@@
Customized Sentence Transformer. Recent advances in sentence transformers Reimers and Gurevych (2019); Cer et al. (2018); Gao et al. (2021) offer state-of-the-art methods to measure semantic relationships between sentences. They utilize a transformer-based architecture combined with pooling and linear layers to produce semantic vector representations of input sentences, enabling the measurement of sentence similarity via cosine similarity. Although traditional sentence transformers cannot be directly applied in our context—since implicit reasoning steps are not explicitly formulated as sentences—they inspire our approach. A naive solution would involve bypassing the tokenization step of sentence transformers and directly feeding implicit or embedded ground-truth reasoning into these models. However, this approach is inadequate, as standard sentence transformers and the LLM we examine typically employ different mappings from tokens to input embeddings, resulting in embedding spaces that differ semantically. To address this issue, we design a customized sentence transformer specifically tailored to handle embeddings from LLMs, thereby accurately measuring the semantic alignment between reasoning.
@@
We design a customized sentence transformer, shown in Fig. 2 (a), denoted as $\mathcal{C}_{\phi}$ (with parameters $\phi$), tailored to comparing semantics between implicit reasoning and ground-truth reasoning. First, we extract the middle five layers of the LLM that we perform CoT on to serve as the backbone of our customized sentence transformer, as these layers have been shown in prior work Liu et al. (2019) to possess optimal language modeling ability and transferability across tasks. Next, we add a pooling layer on top of the transformer layers to aggregate token embeddings from the entire reasoning sequence into a unified vector representation. Finally, we employ a linear layer to project this unified vector into a lower-dimensional semantic embedding space, facilitating more efficient similarity comparisons.
@@
The customized sentence transformer is trained with a specially crafted reasoning pair dataset $\mathcal{G}$, where each data point $(R,S)\in\mathcal{G}$ is a reasoning pair with $R$ being the ground-truth reasoning and $S$ a GPT-4o-mini Menick et al. (2024) generated condensed semantically aligned reasoning of $R$. The prompt to generate $S$ with GPT-4o-mini Menick et al. (2024) is “Please generate the most condensed reasoning text which is semantically aligned with the following reasoning text: RR.” We train the sentence transformer with a contrastive learning Le-Khac et al. (2020) strategy. Specifically, we compute the semantic embeddings of the ground-truth reasoning $\mathcal{C}_{\phi}(\mathcal{T}_{\mathcal{F}}(R_{i}))$ and condensed reasoning $\mathcal{C}_{\phi}(\mathcal{T}_{\mathcal{F}}(S_{i}))$, where $\mathcal{T}_{\mathcal{F}}$ is the mapping from tokens to input embeddings of the LLM. The sentence transformer is trained using contrastive learning to maximize the similarity between embeddings of positive pairs $(R_{i},S_{i})$, while minimizing similarities between all negative pairs $(R_{i},S_{j})$ for $j\neq i$ within the same minibatch. Formally, we learn the sentence transformer with
@@ 3.3 Token-level Efficient Implicit Reasoning Generation
Lightweight Implicit Reasoning Generator. We build the lightweight implicit reasoning generator, shown on the right in Fig. 2 (a), based on an off-the-shelf pruned or distilled LM (e.g., Sheared-LLaMA-1.3B Xia et al. (2024)) from the LLM (e.g., Llama-2-7b-chat-hf Touvron et al. (2023)). We choose to use a pruned or distilled LMs of the original LLMs because existing research Tao et al. (2024); Wang et al. (2025) suggests that those models preserve crucial semantic properties of their original LLMs in their embedding spaces. Then, we incorporate a linear projection layer to map the last-layer hidden embedding generated by the lightweight LM into the embedding space of the LLM to serve as implicit reasoning for the LLM.
@@ 4.1 Experiment Settings
Datasets. We adopt five representative datasets from three different semantic domains used for benchmarking CoT performance. Specifically, we apply mathematical reasoning datasets GSM8K Cobbe et al. (2021), SVAMP Patel et al. (2021), MultiArith ChilleD (2023); Roy and Roth (2015), commonsense reasoning dataset CommonsenseQA Talmor et al. (2019), and symbolic reasoning dataset CoinFlip Krishna et al. (2023). Please see Appendix C for the metadata of the five used datasets.
Examined LLM & Implicit Reasoning Generator. We examine SemCoT on two representative open source LLMs, Llama-2-7b-chat-hf Touvron et al. (2023) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023). For the lightweight LM within the implicit reasoning generator, we utilize the distilled/sheared LMs from their corresponding examined LLMs. Specifically, in the cases that the LLM is Llama-2-7b-chat-hf Touvron et al. (2023), we employ Sheared-LLaMA-1.3B Xia et al. (2024) as the lightweight LM within the implicit reasoning generator. When the backbone LLM is Mistral-7B-Instruct-v0.2 Jiang et al. (2023), we apply mistral-1.1b-testing Optimum Team (2024) accordingly.
Baselines. We adopt state-of-the-art baselines improving CoT efficiency on LLMs. Specifically, (1) Pause Goyal et al. (2024) uses identical implicit reasoning tokens to substitute ground-truth reasoning. Their training strategy includes pretraining and finetuning the LLM for optimal answer accuracy with implicit reasoning. We only implement the finetuning strategy due to the significant cost of LLM pretraining. (2) Progressive encoding methods, such as ICoT-SI Deng et al. (2024) and COCONUT Xu et al. (2025), use implicit tokens to substitute explicit tokens gradually during LLM finetuning. (3) CODI Shen et al. (2025b) adopts a self-distillation strategy, where the teacher and student are both the LLM. The teacher learns to perform explicit reasoning, while the student learns to generate implicit tokens that lead to the correct answer. Simultaneously, both models are trained to align their token embeddings at the final token position of the query. (4) SoftCoT Xu et al. (2025) utilizes a small LM to generate LLM’s hidden reasoning tokens without compressing the number of hidden tokens compared with the original explicit reasoning chain.
@@
Implementation Details. We set the output embedding dimension of the customized sentence transformer to 768. The number of implicit tokens during training is five, and, during evaluation, it is set to one. We optimize both the customized sentence transformer and the implicit reasoning generator with AdamW Loshchilov and Hutter (2019) using the best hyperparameters found. For inference, we allow up to thirty answer tokens to be generated to enforce the LLM to generate the answer instead of the CoT.
@@ 4.2 Effectiveness & Efficiency of SemCoT
In this subsection, we aim to answer RQ1. Specifically, we evaluate our proposed framework SemCoT on two LLMs, Llama-2-7b-chat-hf Touvron et al. (2023) and Mistral-7B-Instruct-v0.2 Jiang et al. (2023). We compare the two LLMs’ answer accuracy and inference time cost when performing CoT with our SemCoT and the state-of-the-art baselines using one implicit reasoning token. We show the results in Table 1. From Table 1, we can make the following observations: (1) from the perspective of effectiveness, our SemCoT achieves the highest answer accuracy compared with the baseline methods in most datasets, showing that our method effectively retains LLM’s reasoning ability in a wide spectrum of NLP tasks and LLM settings. (2) From the perspective of efficiency, SemCoT achieves nearly the fastest implicit reasoning processing time across all datasets and LLMs. In some datasets, SemCoT is slightly slower than SoftCoT, but SemCoT achieves much higher answer accuracy in those cases. In conclusion, SemCoT achieves the optimal performance w.r.t. the trade-off between efficiency and effectiveness.